A viewing screen is separate from double slit source by 1.2m the distance between the two slits is 0.03mm the second order bright fringe (m=2)is 4.5cm from the center line .(a) determine the wavelength of the light.(b) calculate the distance between adjaseet bright fringe



1
Expert's answer
2021-04-13T06:31:55-0400

Given,

Distance between slits and screen "\ D=1.2\ m"

Distance between two slits "d=0.03\ mm"

Second order bight fringe is 4.5 cm from center line

(a) We know,

Position of "2^{nd}" order bright fringe

"y_2=\dfrac{2\lambda D}{d}\\ \Rightarrow \lambda = \dfrac{y_2\times d}{2D}\\ \ \\ \Rightarrow \lambda = \dfrac{4.5\times 3\times 10^{-7}}{2.4}=562.5\ nm"



(b.) The distance between adjacent bright fringe


"\Delta y= \dfrac{\lambda D}{d}=\dfrac{562.5\times 10^{-9}\times1.2}{3\times 10^{-7}}=22.5\times 10^{-3}\ mm"